Blizzard-Wear Glove Company operates a chain of retail kiosks (small retail booths) located at the base of every major ski hill. The kiosks sell ten different styles of ski and mountain biking gloves with identical unit costs and selling prices. A unit is defined as a pair of gloves. Each kiosk has a manager who is paid a fixed salary. Individual salespeople are paid a combination of fixed salary and a sales commission. Blizzard-Wear is trying to determine the desirability of opening another kiosk at Silver Star Mountain in the Okanagan Valley. The new kiosk is expected to have the following revenue and cost relationships

Behaviour

Selling price, per pair of gloves

$  36.00

V

Cost per pair of gloves

23.40

V

Sales commissions per pair of gloves

1.80

V

Kiosk rent

37,500

F

Salaries

145,000

F

Advertising

23,000

F

Other fixed costs

10,500

F

Required (10 marks): (Consider each question independently and use two decimal places.)

  1. What is the annual breakeven point in (a) units sold and (b) revenues? (2 marks)
  2. If the corporate tax rate were 20%, how much additional revenue would have to be earned in order to earn $10,000 after taxes? (2 marks)
  3. If 15,000 units are sold, what will be the store’s operating income (loss)? (2 marks)
  4. If sales commissions were discontinued for individual salespeople in favour of a $10,800 increase in fixed salaries, what would be the annual breakeven point in (a) units sold and (b) revenues?  (2marks)
  5. Refer to the original data. If the store manager were paid $0.36 per unit commission on each unit sold in excess of the breakeven point (as calculated in Part a), what would be the store’s operating income if 35,000 units were sold? (This $0.36 is in addition to both the commission paid to the sales staff and the store manager’s fixed salary based on the original data, not the changes calculated in Part d.) (2 marks)
Answer #1

information given :

selling price per pair of gloves $36
variable cost per pair of gloves :
Cost per pair of gloves $23.40
sales commission $1.80
total variable cost $25.2
fixed cost :
kiosk rent $37500
saaries $145000
advertising $23000
other fixed costs $10500
total fixded cots $216000

a) annual break even points :

(a) break even points (units sold) = fixed costs/contribution per unit

break even points (units sold) = $216000/10.8 = 20000 units

contribution per unit = selling peice per unit - variable cost per unit

contribution per unit = $36 - $25.2 = $10.8

(b)  break even points (revenue) = fixed costs/profit volume ratio

break even points (revenue) = $216000/30% = $720000

profit volume ratio = contribution per unit x 100/ selling price per unit

profit volume ratio = $10.8 x 100/$36 = 30%

b) desired sales to earn $10000 after tax = fixed cost + desired profit or revenue(before tax)/profit volume ratio

desired sales to earn $10000 after tax = ($216000 + $12500)/30% = $761666.66

additional revenue to earn $10000 after tax = desired sales - break even points (revenue)

additional revenue to earn $10000 after tax = $761666.66 - $720000 = $41666.66

desireed profit or revenue (before tax) = desireed profit or revenue (after tax)/(1 - tax rate)

desireed profit or revenue (before tax) = 10000/(1 - 0.20) = 10000/0.80 = $12500

c) If 15,000 units are sold, calculation of the store’s operating income (loss) :

particular amounts in $
sales (15000 x $36) 540000
less : variable costs :
cost (15000 x $23.40) $351000
sales commission (15000 x $1.80) $27000 -378000
fixed costs :
Kiosk rent $37500
salaries $145000
advertising $23000
other fixed costs $10500 -216000
operating loss -54000

d) If sales commissions were discontinued for individual salespeople in favour of a $10800 increase in fixed salaries, calculation of :

annual breakeven point in (a) units sold = fixed costs/contribution per unit

annual breakeven point in (a) units sold = $226800/$12.6 = 18000 units

new fiixed costs = $216000 + $10800 (increase in fixed salaries) = $226800

contribution per unit = selling peice per unit - variable cost per unit

contribution per unit = $36 - $23.40 = $12.6

note - sales commissions were discontinued so not included in variable cost.

and annual breakeven point in (b) revenues = fixed costs/profit volume ratio

annual breakeven point in (b) revenues = $226800/35% = $648000

profit volume ratio = contribution per unit x 100/ selling price per unit

profit volume ratio = $12.6 x 100/$36 = 35%

e) If the store manager were paid $0.36 per unit commission on each unit sold in excess of the breakeven point (as calculated in Part a), calculation of store’s operating income if 35,000 units were sold :

break even point is 20000 units.

particular amounts in $
sales (35000 x $36) 1260000
less : variable costs :
cost (35000 x $23.40) $819000
sales commission (35000 x $1.80) $63000
store manager commision in excess of break even point [(35000-20000) x $0.36] $5400 -887400
fixed costs :
Kiosk rent $37500
salaries $145000
advertising $23000
other fixed costs $10500 -216000
operating income 156600

thank you

please give your feed back and rating